OpenStudy (anonymous):

What is the vector sum of 5 km east and 5 km west?

OpenStudy (anonymous):

i think its 0

OpenStudy (johnweldon1993):

^Indeed  So 0

OpenStudy (anonymous):

thanks to you both. :)

OpenStudy (anonymous):

What is the equation of the line whose y-intercept is 3 and slope is 1? y = x - 3 y = x + 3 y = 3x + 1

OpenStudy (johnweldon1993):

\[\large y = mx + b\] m = slope here b = y-intercept...so if we are given both we have \[\large y = x + 3\]

OpenStudy (anonymous):

thank you, i have another question.

OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (1, 1) and has a y-intercept of 2. x - y + 2 = 0 x + y - 2 = 0 x - y - 2 = 0

OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (1, 1) and has a y-intercept of 2. x - y + 2 = 0 x + y - 2 = 0 x - y - 2 = 0